# R-KV: Redundancy-aware KV Cache Compression for Reasoning Models

### v3

@@ Abstract
Reasoning models have demonstrated impressive performance in self-reflection and chain-of-thought reasoning. However, they often produce excessively long outputs, leading to prohibitively large key-value (KV) caches during inference. While chain-of-thought inference significantly improves performance on complex reasoning tasks, it can also lead to reasoning failures when deployed with existing KV cache compression approaches. To address this, we propose Redundancy-aware KV Cache Compression for Reasoning models (R-KV), a novel method specifically targeting redundant tokens in reasoning models. Our method preserves nearly 100% of the full KV cache performance using only 10% of the KV cache, substantially outperforming existing KV cache baselines, which reaches only 60% of the performance. Remarkably, R-KV even achieves 105% of full KV cache performance with 16% of the KV cache. This KV-cache reduction also leads to a 90% memory saving and a 6.6$\times$ throughput over standard chain-of-thought reasoning inference. Experimental results show that R-KV consistently outperforms existing KV cache compression baselines across two mathematical reasoning datasets.
footnotetext: 🖂 Corresponding to Wen Xiao wxiao@microsoft.com and Junjie Hu junjie.hu@wisc.edu
@@ 1 Introduction
Recent advancements in large language models (LLMs) have demonstrated remarkable capabilities in complex reasoning and self-reflection. However, reasoning models (e.g., DeepSeek-R1 deepseekr1 ) exhibit a critical deployment challenge: their tendency to produce excessively lengthy and redundant reasoning traces results in unsustainable memory demands ConciseReasoning , primarily due to the rapid growth of the key-value (KV) cache during autoregressive generation. For instance, a DeepSeek-R1-Distill-Llama-8B model may generate 32K tokens to solve a complex math problem, consuming 15.5GB of memory to load the model weight and 4.1GB of memory to store the KV cache. This paradigm of long chain-of-thought (CoT) reasoning generation necessitates the development of KV cache compression.
@@
However, existing KV cache compression works snapkv ; h2o ; pyramidkv ; headkv ; streamingllm primarily handle long input prompts but do not explore extensively for long generation outputs. Furthermore, based on our observation (§2.2), standard KV-cache compression methods that rely on simple attention-based importance filtering often fail because the repetitive sections generate high attention signals for themselves. Naively pruning tokens with “low attention weight” may remove crucial but scattered bits of reasoning, or over-retain duplicative self-reflections that appear to have high attention. This observation motivates our exploration of redundancy-aware compression strategies, which selectively retain “important and non-repetitive context” during decoding to preserve the model’s critical reasoning ability.
@@
Figure 2: Comparison of generation length and average 1-/2-gram frequency for reasoning models and ground truth of MATH-500 hendrycks2021measuring and AIME 2024 AIME2024 . . Reasoning models generate substantially longer responses with 8-14$\times$ more tokens, and show higher word repetition with 5-7$\times$ higher frequency.
@@ 2.1 Redundancy in Reasoning Models
As noted in ConciseReasoning , reasoning models often generate a detailed chain of thoughts and multiple reflection steps, resulting in significantly longer responses than standard models. Figure 2 shows that both reasoning models (i.e., DeepSeek-R1-Distill-Llama-8B, DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Qwen-14B) generate more than 8$\times$ longer generation output compared to the ground truth on two popular math reasoning datasets. However, not all of the additional tokens contribute meaningful content, as much of the decoded context is dominated by repetition. Figure 2 also shows that the average frequency of 1- to 2-grams is consistently higher in the generation output of reasoning models than in ground truth, indicating greater repetitions in the generated outputs of reasoning models.
@@ 2.2 Failure of Existing KV Compression Methods to Handle Redundancy
Most existing KV cache compression methods prioritize token selection based primarily on tokens’ contextual importance, typically measured through attention scores between key and query tokens snapkv ; h2o . While this approach effectively retains critical context, it fails to account for redundancy—particularly problematic in reasoning models. In such models, we find that repetitive content often receives disproportionately high attention scores, as it closely mirrors previously generated repetitive text. As a result, redundant tokens are excessively retained, unnecessarily inflating the KV cache size without providing additional meaningful new information. In Figure 3, we visualize the cached tokens (inside red boxes) selected by a popular attention-based KV cache method (i.e., SnapKV), showing many repetitions related to self-reflection and conclusion to the final answer.
@@ 3.1 Decoding-time Compression
Different from existing KV cache compression methodssnapkv ; pyramidkv ; h2o that focus on the prefilling stage to manage long-context inputs, our R-KV focuses on the decoding stage for reasoning models—a distinctive setting where the generated output is significantly longer than the prompt.
Specifically, R-KV allocates memory for two components: a cache of budget size $B_{\text{budget}}$ to store retained KV tokens, and a buffer of size $B_{\text{buffer}}$ for newly generated text tokens. The total memory requirement is thus $B_{\text{total}}=B_{\text{budget}}+B_{\text{buffer}}$. After the model generates each fixed-length text segment in the buffer, R-KV performs KV cache compression. At the end of each text segment, the last $\alpha$ tokens are always retained in the cache as observation tokens, following prior work snapkv . Next, we concatenate the existing $B_{\text{budget}}$ tokens in the cache with the first $B_{\text{buffer}}-\alpha$ tokens in the buffer, resulting in $n=B_{\text{budget}}+B_{\text{buffer}}-\alpha$ candidate KV tokens. Each candidate is assigned a selection score (§3.4), and we select the top $k=B_{\text{budget}}-\alpha$ tokens to fit in the rest of the cache budget, in addition to the $\alpha$ observation tokens. This process compresses the KV cache while preserving critical context, enabling efficient memory utilization during autoregressive decoding.
@@ 3.2 Importance Scoring via Attention Weights
Following attention-based methods (e.g., SnapKV snapkv , PyramidKV pyramidkv ), R-KV estimates token importance using attention weights, leveraging the intuition that tokens receiving higher attention contribute more to decoding and are thus more critical for preserving model performance. Specifically, we compute each key token’s attention scores received from the last $\alpha$ observation tokens during decoding. In addition to the standard multi-head attention mainly adopted by the prior works snapkv , we also propose the importance score estimation using the grouped-query attention. Below, we detail the estimation on top of these two popular attention mechanisms used by current LLMs.
@@
We use ${\bm{A}}^{h}$ hereafter to denote the attention weights calculated using either MHA or GQA. Note that the per-token importance scores derived from ${\bm{A}}^{h}$ may contain outliers with excessively high values, resulting in unstable estimation of importance scores. To mitigate this influence, we follow the prior work snapkv and apply a max-pooling operation to these per-token importance scores over a sliding window of size $2W$ across recent tokens. Specifically, we denote $A^{h}_{j,i}$ as the attention score from the $j$-th query to the $i$-th key in ${\bm{A}}^{h}$. We obtain the stabilized attention score $\tilde{{\bm{A}}}^{h}$ by computing its $(i,j)$ entry, and finally obtain the importance score of retaining the $i$-th token in the KV cache as $\text{I}^{h}_{i}$ for each attention head $h$, as shown below:
@@ 4.1 Experimental Setup
In our experiments, we use variants of the DeepSeek-R1 distilled model: DeepSeek-R1-Distill-Llama-8B, and DeepSeek-R1-Distill-Qwen-14B deepseekr1 , which we refer to as R1-Llama-8B and R1-Qwen-14B, respectively, for brevity throughout the paper.
We evaluate the models’ mathematical reasoning capabilities using three benchmarks: MATH-500 hendrycks2021measuring and AIME 2024 AIME2024 .
@@
We compare our method against SnapKV snapkv , originally designed for long prefilling. To adapt it for decoding, we apply the same compression interval as our method, i.e., compressing the KV cache every 128 decoding steps using identical $B_{\text{budget}}$ and $B_{\text{buffer}}$. Our approach focuses on improving KV cache eviction through a hybrid strategy, and we therefore restrict comparison to state-of-the-art attention-based eviction methods. Budget allocation techniques (e.g., head-level headkv and layer-level pyramidkv ) are orthogonal to our work and not included. We also report results for FullKV, which retains the full KV cache and serves as the gold standard for decoding quality.
@@
[…]
@@ 4.2 Results
The accuracy performance of R-KV compared with all baselines is shown in Figure 4, with detailed accuracy numbers in Section B.2. The KV cache budget ratio is calculated based on the KV cache budget and the average generation length of tokens, i.e., R1-Llama-8B: $2,979.1$ on MATH-500 and $15,535.8$ on AIME24; R1-Qwen-14B: $2,833.04$ on MATH-500 and $12,402$ on AIME24. Our method significantly outperforms the baseline SnapKV, achieving up to $40\%$ Acc. improvement. We provide two KV cache budget and performance analysis. Fixed budget analysis is more practical because when the model outputs longer (i.e., from $2,979.1$ on MATH-500 to $15,535.8$ on AIME24), the KV cache budget needed for lossless compression increases less (i.e., $512$). In the KV cache budget ratio perspective, the changes of lossless compression ratio is dominated by generation length.
@@ 5.1 How to Choose $\lambda$?
Figure 6 shows the distributions of the Importance Score ($\text{I}^{h}$) and Redundancy Estimation ($\text{R}^{h}$) for head $h=0$ at the top layer ($N_{\text{layer}}=31$). The figure reveals that $\text{I}^{h}$ is sparse and dominated by a few outlier values, while the similarity distributions (which inform $\text{R}^{h}$) are relatively dense. When $\lambda=0$, the token retention strategy is overned entirely by Redundancy Estimation ($\text{R}^{h}$). As shown in Figure 6, the initial four tokens are not guaranteed to be preserved. As highlighted by prior work streamingllm , evicting these initial tokens can severely impair the generative capabilities of LLMs. Therefore, it is crucial to select a $\lambda$ value that starts from at least $0.01$. On the other hand, as $\lambda$ increases beyond $0.1$, the selection metric becomes increasingly dominated by attention scores. These observations suggest that an optimal $\lambda$ lies within the range of $0.01\leq\lambda\leq 0.1$, effectively balancing the contributions of Importance Score and Redundancy Estimation.
@@ 5.3 Efficiency Analysis
Table 1: Memory saving, throughput, and decoding-time comparison for Llama3-8B under various generation length and KV cache compression budget settings.
@@
We evaluate end-to-end throughput under both ratio-based and fixed KV cache budgets. R-KV consistently enables much larger batch sizes and higher throughput than FullKV, with benefits becoming more pronounced at longer sequence lengths. For example, at a sequence length of 16K, R-KV achieves up to 9$\times$ larger batch sizes and over 6.6$\times$ higher throughput under a 10% compression ratio, and 13.4$\times$ larger batch sizes with 9.2$\times$ throughput under a fixed budget of 1024. Detailed analysis are provided in Appendix C.2.
@@ 6 Related Work
The optimization of KV cache memory efficiency in LLMs has garnered increasing attention as model sizes and context windows expand. Existing approaches primarily fall into three categories: dynamic token evictionsnapkv ; fastgen ; scissorhands , quantizationhooper2024kvquant ; liu2024kivi ; yue2024wkvquantquantizingweightkeyvalue , mergingzhang2024cam ; kim2023compressed ; nawrot2024dynamic , and low-rank decompositionsun2025shadowkvkvcacheshadows ; saxena2024eigenattentionattentionlowrank ; zhang2024lorclowrankcompressionllms . Previous eviction methods like SnapKVsnapkv , PyramidKVpyramidkv , Ada-KVadakv , HeadKVheadkv dynamically prune tokens based on attention scores, but mainly focus on evicting tokens for prefilling stage. StreamingLLMstreamingllm and H2Oh2o are proposed for decoding. However, these general-purpose techniques often struggle with reasoning-intensive tasks, where aggressive eviction risks disrupting critical intermediate steps in CoT, and suffers from reasoning models’ inherent redundency.
@@
Recent works in efficient reasoning focus on training the model to generate less CoT without sacrificing performance. li2025adaptivegrouppolicyoptimization ; yang2025thinkneedselfadaptivechainofthought ; kimiteam2025kimik15scalingreinforcement use RL optimization with length penalty rewards to encourage models to produce more concise chains-of-thought (CoT). cui2025stepwiseperplexityguidedrefinementefficient ; han2025tokenbudgetawarellmreasoning employs variable-length CoT datasets to supervised fine-tune (SFT) the LLM to reduce token usage while preserving reasoning correctness. Both RL and SFT methods require additional training. han2025tokenbudgetawarellmreasoning ; liu2025thoughtmanipulationexternalthought ; ma2025reasoningmodelseffectivethinking use test-time prompting to reduce generation length, but these methods may hurt the performance. As a KV cache compression work for reasoning models, R-KV is able to achieve lossless compression without extensive training and prompting.
@@ A.1 Algorithm
The pseudo-code of the method is shown in Algorithm 1.
@@
1:procedure R-KV($({\bm{K}}_{\text{full}},{\bm{V}}_{\text{full}}),L_{\text{full}},L_{\text{budget}},Q_{\text{obs}},\alpha,B_{\text{budget}},B_{\text{buffer}},T,\beta,\lambda,\epsilon,H,d_{k}$)
@@
7:     $N_{c}\leftarrow L_{\text{full}}-\alpha$triangleright Number of candidate tokens
@@
9:         return $({\bm{K}}_{\text{full}},{\bm{V}}_{\text{full}})$triangleright Not enough candidates to prune beyond budget
@@
12:         Compute attention matrix ${\bm{A}}^{h}\in\mathbb{R}^{\alpha\times N_{c}}$ using $Q_{\text{obs}}^{h}$ and ${\bm{K}}_{\text{cand}}^{h}$triangleright Handles MHA/GQA as per Eqs. (1)-(3) from text
@@
14:              $I^{\prime}_{k,h}\leftarrow\frac{1}{\alpha}\sum_{q=0}^{\alpha-1}({\bm{A}}^{h})_{qk}$triangleright $q$: observation token, $k$: candidate token
@@
20:         ${\bm{S}}^{h}\leftarrow{\bm{K}}_{\text{norm}}^{h}({\bm{K}}_{\text{norm}}^{h})^{\top}$triangleright Cosine Similarity Matrix Computation, similarity matrix ${\bm{S}}^{h}\in\mathbb{R}^{N_{c}\times N_{c}}$
@@
24:         $B_{uv}^{h}\leftarrow(({\bm{S}}^{h})_{uv}>T?1:0)$ for $u,v\in\{0,\dots,N_{c}-1\}$triangleright Identify Highly Similar Pairs
@@
27:              $T_{u,\beta}^{h}\leftarrow\text{subset of }T_{u}^{h}\text{ with up to }\beta\text{ largest indices }v$.
@@
44:         $\text{AggScore}_{k}\leftarrow\text{mean}_{h}(\text{Score}_{k,h})$triangleright Aggregate scores across heads
@@
46:     $Idx_{sel}\leftarrow\text{indices of top-}B_{budget}\text{ tokens from }\{0,\dots,N_{c}-1\}\text{ based on }\text{AggScore}$
@@
48:     ${\bm{K}}_{comp}\leftarrow\text{concatenate}({\bm{K}}_{cand\_sel},{\bm{K}}_{obs})$triangleright Order might vary
@@
50:     $L_{prev\_comp}\leftarrow B_{budget}+\alpha$triangleright Update length for next cycle

### v4

@@ Abstract
[…]
footnotetext: 🖂 Corresponding to Zefan Cai zefncai@gmail.com, Wen Xiao wxiao@microsoft.com and Junjie Hu junjie.hu@wisc.edu
@@ 1 Introduction
Recent advancements in large language models (LLMs) have demonstrated remarkable capabilities in complex reasoning and self-reflection. However, reasoning models (e.g., DeepSeek-R1 DeepSeek-AI et al. (2025)) exhibit a critical deployment challenge: their tendency to produce excessively lengthy and redundant reasoning traces results in unsustainable memory demands Fatemi et al. (2025), primarily due to the rapid growth of the key-value (KV) cache during autoregressive generation. For instance, a DeepSeek-R1-Distill-Llama-8B model may generate 32K tokens to solve a complex math problem, consuming 15.5GB of memory to load the model weight and 4.1GB of memory to store the KV cache. This paradigm of long chain-of-thought (CoT) reasoning generation necessitates the development of KV cache compression.
@@
However, existing KV cache compression works Li et al. (2024); Zhang et al. (2023); Cai et al. (2024); Fu et al. (2024); Xiao et al. (2024) primarily handle long input prompts but do not explore extensively for long generation outputs. Furthermore, based on our observation (§2.2), standard KV‑cache compression methods that rely on simple attention‑based importance filtering often fail because the repetitive sections generate high attention signals for themselves. Naively pruning tokens with “low attention weight” may remove crucial but scattered bits of reasoning, or over‑retain duplicative self‑reflections that appear to have high attention. This observation motivates our exploration of redundancy-aware compression strategies, which selectively retain “important and non-repetitive context” during decoding to preserve the model’s critical reasoning ability.
@@
[…]
@@ 2.1 Redundancy in Reasoning Models
[…]
@@ 2.2 Failure of Existing KV Compression Methods to Handle Redundancy
Most existing KV cache compression methods prioritize token selection based primarily on tokens’ contextual importance, typically measured through attention scores between key and query tokens Li et al. (2024); Zhang et al. (2023). While this approach effectively retains critical context, it fails to account for redundancy—particularly problematic in reasoning models. In such models, we find that repetitive content often receives disproportionately high attention scores, as it closely mirrors previously generated repetitive text. As a result, redundant tokens are excessively retained, unnecessarily inflating the KV cache size without providing additional meaningful new information. In Figure 3, we visualize the cached tokens (inside red boxes) selected by a popular attention-based KV cache method (i.e., SnapKV), showing many repetitions related to self-reflection and conclusion to the final answer.
@@ 3.1 Decoding-time Compression
Different from existing KV cache compression methodsLi et al. (2024); Cai et al. (2024); Zhang et al. (2023) that focus on the prefilling stage to manage long-context inputs, our R-KV focuses on the decoding stage for reasoning models—a distinctive setting where the generated output is significantly longer than the prompt.
Specifically, R-KV allocates memory for two components: a cache of budget size $B_{\text{budget}}$ to store retained KV tokens, and a buffer of size $B_{\text{buffer}}$ for newly generated text tokens. The total memory requirement is thus $B_{\text{total}}=B_{\text{budget}}+B_{\text{buffer}}$. After the model generates each fixed-length text segment in the buffer, R-KV performs KV cache compression. At the end of each text segment, the last $\alpha$ tokens are always retained in the cache as observation tokens, following prior work Li et al. (2024). Next, we concatenate the existing $B_{\text{budget}}$ tokens in the cache with the first $B_{\text{buffer}}-\alpha$ tokens in the buffer, resulting in $n=B_{\text{budget}}+B_{\text{buffer}}-\alpha$ candidate KV tokens. Each candidate is assigned a selection score (§3.4), and we select the top $k=B_{\text{budget}}-\alpha$ tokens to fit in the rest of the cache budget, in addition to the $\alpha$ observation tokens. This process compresses the KV cache while preserving critical context, enabling efficient memory utilization during autoregressive decoding.
@@ 3.2 Importance Scoring via Attention Weights
Following attention-based methods (e.g., SnapKV Li et al. (2024), PyramidKV Cai et al. (2024)), R-KV estimates token importance using attention weights, leveraging the intuition that tokens receiving higher attention contribute more to decoding and are thus more critical for preserving model performance. Specifically, we compute each key token’s attention scores received from the last $\alpha$ observation tokens during decoding. In addition to the standard multi-head attention mainly adopted by the prior works Li et al. (2024), we also propose the importance score estimation using the grouped-query attention. Below, we detail the estimation on top of these two popular attention mechanisms used by current LLMs.
@@
We use ${\bm{A}}^{h}$ hereafter to denote the attention weights calculated using either MHA or GQA. Note that the per-token importance scores derived from ${\bm{A}}^{h}$ may contain outliers with excessively high values, resulting in unstable estimation of importance scores. To mitigate this influence, we follow the prior work Li et al. (2024) and apply a max-pooling operation to these per-token importance scores over a sliding window of size $2W$ across recent tokens. Specifically, we denote $A^{h}_{j,i}$ as the attention score from the $j$-th query to the $i$-th key in ${\bm{A}}^{h}$. We obtain the stabilized attention score $\tilde{{\bm{A}}}^{h}$ by computing its $(i,j)$ entry, and finally obtain the importance score of retaining the $i$-th token in the KV cache as $\text{I}^{h}_{i}$ for each attention head $h$, as shown below:
@@ 4.1 Experimental Setup
In our experiments, we use variants of the DeepSeek-R1 distilled model: DeepSeek-R1-Distill-Llama-8B, and DeepSeek-R1-Distill-Qwen-14B DeepSeek-AI et al. (2025), which we refer to as R1-Llama-8B and R1-Qwen-14B, respectively, for brevity throughout the paper.
We evaluate the models’ mathematical reasoning capabilities using three benchmarks: MATH-500 Hendrycks et al. (2021) and AIME 2024 MAA (2024).
@@
We compare our method against SnapKV Li et al. (2024), originally designed for long prefilling. To adapt it for decoding, we apply the same compression interval as our method, i.e., compressing the KV cache every 128 decoding steps using identical $B_{\text{budget}}$ and $B_{\text{buffer}}$. Our approach focuses on improving KV cache eviction through a hybrid strategy, and we therefore restrict comparison to state-of-the-art attention-based eviction methods. Budget allocation techniques (e.g., head-level Fu et al. (2024) and layer-level Cai et al. (2024)) are orthogonal to our work and not included. We also report results for FullKV, which retains the full KV cache and serves as the gold standard for decoding quality.
@@
We set the maximum generation length to 16,384 tokens for MATH-500 and 32,768 tokens for AIME 2024 and AIME 2025, because further increasing the generation length has shown no improvement on model performance on these datasets from our attempts. We find that using greedy decoding to evaluate long-output reasoning models results in significant variability across different setups. Following existing works DeepSeek-AI et al. (2025), we utilize pass@$k$ evaluation (Chen et al., 2021) and report pass@1 using a non-zero temperature. We use the recommended sampling temperature and top-$p$ value for each model, i.e., sampling temperature of $0.6$ and a top-$p$ value of $0.95$ for DeepSeek-R1 Distilled models. We generate $64$ responses for each question. Pass@1 is then calculated as $\text{Pass@1}=\frac{1}{k}\sum_{i=1}^{k}p_{i}$, where $p_{i}$ denotes the correctness of the $i$-th response. This method provides more reliable performance estimates.
@@ 4.2 Results
The accuracy performance of R-KV compared with all baselines is shown in Figure 4, with detailed accuracy numbers in Section˜B.2. The KV cache budget ratio is calculated based on the KV cache budget and the average generation length of tokens, i.e., R1-Llama-8B: $2,979.1$ on MATH-500 and $15,535.8$ on AIME24; R1-Qwen-14B: $2,833.04$ on MATH-500 and $12,402$ on AIME24. Our method significantly outperforms the baseline SnapKV, achieving up to $40\%$ Acc. improvement. We provide two KV cache budget and performance analysis. Fixed budget analysis is more practical because when the model outputs longer (i.e., from $2,979.1$ on MATH-500 to $15,535.8$ on AIME24), the KV cache budget needed for lossless compression increases less (i.e., $512$). In the KV cache budget ratio perspective, the changes of lossless compression ratio is dominated by generation length.
@@ 5.1 How to Choose $\lambda$?
Figure 6 shows the distributions of the Importance Score ($\text{I}^{h}$) and Redundancy Estimation ($\text{R}^{h}$) for head $h=0$ at the top layer ($N_{\text{layer}}=31$). The figure reveals that $\text{I}^{h}$ is sparse and dominated by a few outlier values, while the similarity distributions (which inform $\text{R}^{h}$) are relatively dense. When $\lambda=0$, the token retention strategy is overned entirely by Redundancy Estimation ($\text{R}^{h}$). As shown in Figure 6, the initial four tokens are not guaranteed to be preserved. As highlighted by prior work Xiao et al. (2024), evicting these initial tokens can severely impair the generative capabilities of LLMs. Therefore, it is crucial to select a $\lambda$ value that starts from at least $0.01$. On the other hand, as $\lambda$ increases beyond $0.1$, the selection metric becomes increasingly dominated by attention scores. These observations suggest that an optimal $\lambda$ lies within the range of $0.01\leq\lambda\leq 0.1$, effectively balancing the contributions of Importance Score and Redundancy Estimation.
@@ 5.3 Efficiency Analysis
Table 1: Memory saving, throughput, and decoding‐time comparison for Llama3-8B under various generation length and KV cache compression budget settings.
@@
[…]
@@ 6 Related Work
The optimization of KV cache memory efficiency in LLMs has garnered increasing attention as model sizes and context windows expand. Existing approaches primarily fall into three categories: dynamic token evictionLi et al. (2024); Ge et al. (2023); Liu et al. (2023), quantizationHooper et al. (2024); Liu et al. (2024); Yue et al. (2024), mergingZhang et al. (2024b); Kim et al. (2023); Nawrot et al. (2024), and low-rank decompositionSun et al. (2025); Saxena et al. (2024); Zhang et al. (2024a). Previous eviction methods like SnapKVLi et al. (2024), PyramidKVCai et al. (2024), Ada-KVFeng et al. (2024), HeadKVFu et al. (2024) dynamically prune tokens based on attention scores, but mainly focus on evicting tokens for prefilling stage. StreamingLLMXiao et al. (2024) and H2OZhang et al. (2023) are proposed for decoding. However, these general-purpose techniques often struggle with reasoning-intensive tasks, where aggressive eviction risks disrupting critical intermediate steps in CoT, and suffers from reasoning models’ inherent redundency.
@@
Recent works in efficient reasoning focus on training the model to generate less CoT without sacrificing performance. Li et al. (2025); Yang et al. (2025); Team et al. (2025) use RL optimization with length penalty rewards to encourage models to produce more concise chains-of-thought (CoT). Cui et al. (2025); Han et al. (2025) employs variable-length CoT datasets to supervised fine-tune (SFT) the LLM to reduce token usage while preserving reasoning correctness. Both RL and SFT methods require additional training. Han et al. (2025); Liu et al. (2025); Ma et al. (2025) use test-time prompting to reduce generation length, but these methods may hurt the performance. As a KV cache compression work for reasoning models, R-KV is able to achieve lossless compression without extensive training and prompting.
@@ 7 Conclusion
8 Acknowledgement
Research reported in this publication was partially supported by the National Science Foundation under Award Number IIS-2449768. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Science Foundation.
@@ A.1 Algorithm
The pseudo-code of the method is shown in Algorithm˜1.
@@
[…]
@@
7:  $N_{c}\leftarrow L_{\text{full}}-\alpha$ $\triangleright$ Number of candidate tokens
@@
9:   return $({\bm{K}}_{\text{full}},{\bm{V}}_{\text{full}})$ $\triangleright$ Not enough candidates to prune beyond budget
@@
12:   Compute attention matrix ${\bm{A}}^{h}\in\mathbb{R}^{\alpha\times N_{c}}$ using $Q_{\text{obs}}^{h}$ and ${\bm{K}}_{\text{cand}}^{h}$ $\triangleright$ Handles MHA/GQA as per Eqs. (1)-(3) from text
@@
14:     $I^{\prime}_{k,h}\leftarrow\frac{1}{\alpha}\sum_{q=0}^{\alpha-1}({\bm{A}}^{h})_{qk}$ $\triangleright$ $q$: observation token, $k$: candidate token
@@
20:   ${\bm{S}}^{h}\leftarrow{\bm{K}}_{\text{norm}}^{h}({\bm{K}}_{\text{norm}}^{h})^{\top}$ $\triangleright$ Cosine Similarity Matrix Computation, similarity matrix ${\bm{S}}^{h}\in\mathbb{R}^{N_{c}\times N_{c}}$
@@
24:   $B_{uv}^{h}\leftarrow(({\bm{S}}^{h})_{uv}>T?1:0)$ for $u,v\in\{0,\dots,N_{c}-1\}$ $\triangleright$ Identify Highly Similar Pairs
@@
[…]
@@
44:   $\text{AggScore}_{k}\leftarrow\text{mean}_{h}(\text{Score}_{k,h})$ $\triangleright$ Aggregate scores across heads
@@
[…]
@@
48:  ${\bm{K}}_{comp}\leftarrow\text{concatenate}({\bm{K}}_{cand\_sel},{\bm{K}}_{obs})$ $\triangleright$ Order might vary
@@
50:  $L_{prev\_comp}\leftarrow B_{budget}+\alpha$ $\triangleright$ Update length for next cycle